OpenStudy (anonymous):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for two different years. Number of years 1 2 3 Option 1 (amount in dollars) 1300 1600 1900 Option 2 (amount in dollars) 1120 1254.40 1404.93 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C: Be

OpenStudy (anonymous):

So far, I wrote: Part A: Option one can be a linear as it makes a straight line when graphed. Option two, however, is exponential, because it does not make a straight line. I don't quite understand how to do this problem, could someone help explain it to me?

OpenStudy (solomonzelman):

I never liked long and annoying word problems like this.

OpenStudy (wolf1728):

It's a compound interest problem and I like those.

OpenStudy (anonymous):

I really just need help on part b, I've been stuck on it for a while.

OpenStudy (wolf1728):

Part A - Option 1 is linear and it is simple interest

OpenStudy (wolf1728):

Okay on to part B

OpenStudy (wolf1728):

Option 1 Total = 1,000 + (1 + n*.3) n = number of years the interest rate is a whopping 30% per year (but it is simple interest). Here's some calculations from year zero to year 21 0 1,000.00 1 1,300.00 2 1,600.00 3 1,900.00 4 2,200.00 5 2,500.00 6 2,800.00 7 3,100.00 8 3,400.00 9 3,700.00 10 4,000.00 11 4,300.00 12 4,600.00 13 4,900.00 14 5,200.00 15 5,500.00 16 5,800.00 17 6,100.00 18 6,400.00 19 6,700.00 20 7,000.00 21 7,300.00

OpenStudy (wolf1728):

Option 2 Total = $1,000 * (1 + rate)^n The rate here is 12% but it is compound interest and again 'n' is years Here are some calculations from year zero to year 21. 0 1,000.00 1 1,120.00 2 1,254.40 3 1,404.93 4 1,573.52 5 1,762.34 6 1,973.82 7 2,210.68 8 2,475.96 9 2,773.08 10 3,105.85 11 3,478.55 12 3,895.98 13 4,363.49 14 4,887.11 15 5,473.57 16 6,130.39 17 6,866.04 18 7,689.97 19 8,612.76 20 9,646.29 21 10,803.85 Note that at year 16, option 2 produces more money than option 1. Although option 2 has a much smaller interest rate (12%) (compared to 30%), it is compound interest.

OpenStudy (solomonzelman):

Crunchberries ni _ _ er damn! But damn!!!!!!!!!!!!!!!

OpenStudy (wolf1728):

and dat's da name of dat tune !!!

OpenStudy (anonymous):

What is the rate number? 0.12 didn't work

OpenStudy (wolf1728):

option 1 rate is 30% but use .3 option 2 is 12% but use .12

OpenStudy (anonymous):

Oh! I get it now!! Thank you so so much! You really helped me out a lot!

OpenStudy (wolf1728):

Let's try 5 years at option 2 total = $1,000 * (1.12)^5 total = $1,000 * 1.76234168 total = $1,762.34

OpenStudy (wolf1728):

okay momoluvgd glad I helped you out (I had a feeling I'd be good at this compound interest stuff)